I am desperate for some HELP!
Determine whether each of the distributions given below represents a probability distribution.
Justify your answers.
(A)
x_____|___1_____2_____3_______4
P (x) …| 1/5 ….5/12….1/3………1/6
(B)
X____| ___3_____6________8
P (x) | …0.2….2/5…………0.3
(C)
X____|___20_____35____40____50
P(x) | …..0.3…….-0.1…..0.5…….0.3

Answer by ewatrrr(24785)


 
Hi,
the idea is too check to see if:
1) the probabilites in @ case total 1
2) there is no negative probabilities listed as an option
A 1/5 + 5/12 + 1/3 + 1/6 = 12/60 + 25/60 + 20/60 + 10/60 = 67/60 ≠ 1
B 0.2 + 2/5 + 0.3 = 2/10 + 4/10 + 3/10 = 9/10 ≠ 1
C P(x)is -0.1 not possible... cannot have a 'negative' probability